OpenStudy (anonymous):

An entrepeneur estimates that total profit( profit = total cost) from his proposed company will be given the function P(x)=x^3-4x^2+3x-12, when P is in hundreds of dollars and x is the number of years elapsed after the company starts operations. in how many years (x) will the company break even (no profit, no loss)? A)1year B)2 years C)3 years D)4 years E)6 years

OpenStudy (anonymous):

Hmm, did you try plugging numbers in? :).. such as the years?

OpenStudy (anonymous):

Try plugging these in, and simplify it out and check if y = 0

OpenStudy (anonymous):

if y = 0 that's your break even point.

OpenStudy (anonymous):

yes and i still dont under stand it

OpenStudy (anonymous):

what do you not understand?

OpenStudy (anonymous):

how to do this type of work

OpenStudy (anonymous):

x stands for time, and in this case, x is the number of years- so you CHECK the amount of money the company makes in the certain number of years by multiplying the number of years (x) into the equation. Okay I will walk you through it. Do you know how to multiply exponents?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so would i do P(x)=1^3-4*1^2+3*1-12

OpenStudy (anonymous):

yeah thats right

OpenStudy (anonymous):

P(x)=1-4+3-12 P(x)=3+3-12 P(x)=6-12 P(x)=6

OpenStudy (anonymous):

don't forget about the signs.. it supposed to be -12

OpenStudy (anonymous):

P(x)=2^3-4*2^2+3*2-12 P(x)=8-16+6-12 P(x)=8+6-12 P(x)=14-12 P(x)=2

OpenStudy (anonymous):

hmm..it's -14

OpenStudy (anonymous):

just do it from left to right when doing your addition to subtraction. After you got the exponents out- it's just left to right with subtraction and addition.

OpenStudy (anonymous):

that is what i did

OpenStudy (anonymous):

So your steps were perfect, just do literally: 8-16 = -8 +6 = -2 - 12 = -14

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

good job, keep going!

OpenStudy (anonymous):

How about 4? Because if you plug that in, y should equal to zero :)

OpenStudy (anonymous):

ok it is not A, B, or C

OpenStudy (anonymous):

x^3-4x^2+3x-12 X^3= 4^3 = 64 X^3(4) = 4^2 =16(4) = 64 and you have to minus so this equals to 0 and the next part.. 3(x) = 3(4) = 12 - 12 = 0 So that is when they are at the break even point. :) D)4 years

OpenStudy (anonymous):

but you understand now right? ;p

OpenStudy (anonymous):

ok thank you yes i understand now

OpenStudy (anonymous):

that's good, good luck with your maths.